An electric car is being designed to have an average power output of 4,200 W for 2 h before needing to be recharged. (Assume there is no wasted energy.) (a) How much energy would be stored in the charged batteries?

____J

(b) The batteries operate on 50 V. What would the current be when they are operating at 4,200 W?

_____A

(c) To be able to recharge the batteries in 1 h, how much power would have to be supplied to them?

_____W
Answer #1

answer a)

Power = Work/ change in time

work(energy)= 4200*2*3600=   30240000 j

B) ipv.gif

I=4200/50 =84 amp.

C)   

joules = watts
Answer #2

a) E =PT

E = 4200 *(2*60*60)

= 30240 KJ

B) P=VI

4200 = 50 I

I = 84 AMPERE

C) 4200W REQUIRED TO CHARGE IN 2 HR

8400 W REQUIRED TO CHARGE IN 1HR
Answer #3

A)


power ouput = energy/time...

energy E = Power*time = 4200*2*3600 = 3.024*10^7 J...

B) P = V*I..
current I = P/V = 4200/50 = 84 A..

C)P1*t1 = P2*t2...

P2 = P1*t1/t2 = 4200*2 = 8400W
Answer #4

Power=Work Done/Time

Energy=Power*Time

           4200*2*3600

          =30.24MJ

power=IV*i

4200=V*I

I=84A

power=Work Done /Time

         = V*I/T

         =8.4KW